OpenStudy (anonymous):

conad Rationalize denominator and simplify. A. 2/ sqrt 7+ sqrt 5 B. 6/sqrt 27

OpenStudy (anonymous):

\[\frac{ 2 }{ \sqrt{7}+\sqrt{5} } \] ?

OpenStudy (anonymous):

is that what you want to rationalize?

OpenStudy (luigi0210):

Multiplying by the conjugate usually works.

OpenStudy (anonymous):

I was just helping you on your exact same post...

OpenStudy (anonymous):

No one is going to just give you the answer, we will help you learn but you're the one who has to put the effort in finding the result.

OpenStudy (anonymous):

So once again, \[\frac{ 2 }{ \sqrt{7}+\sqrt{5} } \times \frac{ \sqrt{7}-\sqrt{5} }{ \sqrt{7}-\sqrt{5} }\] what do you get?